OpenStudy (kenkaneki12280000):

The coordinates below represent two linear equations. How many solutions does this system of equations have? Line 1 x y –3 5 0 4 Line 2 x y 0 4 6 2 A. 0 B. exactly 1 C. exactly 2 D. infinitely many @jameshorton

OpenStudy (kenkaneki12280000):

@jameshorton

OpenStudy (jameshorton):

d

OpenStudy (kenkaneki12280000):

ready for the next one

OpenStudy (kenkaneki12280000):

The coordinates below represent two linear equations. How many solutions does this system of equations have? Line 1 x y –4 8 4 6 Line 2 x y –1 1 3 5 A. 0 B. exactly 1 C. exactly 2 D. infinitely many

OpenStudy (jameshorton):

0

OpenStudy (kenkaneki12280000):

QUESTION 3 srry caps were on Complete the equations of the system in slope-intercept form. Use a decimal for the slope if necessary. Line 1 x y 0 3 1 5 Line 2 x y –1 1 –2 –1 Line 1: y = x + Line 2: y = x +

OpenStudy (jameshorton):

y=2x+3 line 1

OpenStudy (jameshorton):

line 2 same as line 1

OpenStudy (jameshorton):

i said a the first one was a

OpenStudy (kenkaneki12280000):

yea noo i meant the one we just answered

OpenStudy (kenkaneki12280000):

What is the solution to the system? Line 1 x y –1 –2 1 2 Line 2 x y –2 10 4 1 ( , )

OpenStudy (kenkaneki12280000):

@jameshorton

OpenStudy (jameshorton):

Point slope form again -3-9/3+1= -3 y+1=-3(x-9) y+1=-3x+27 y=-3x+26 5+1/2+1=2 y+1=2(x+1) y+1=2x+2 y=2x+1 For your system of equations you have y={y=-3x+26} y={y=2x+1} There is a couple ways you can solve this. You could either use a subtraction method, or a substitution method. since they are both y = something i would just substitute one y in one or the other equation. 2x+1=-3x+26 5x+1=26 5x=25 x=5 So now you just plug in x = 5 for each quotation and you have your y's for both solution points y=-3(5)+26 y=11 y=2(5)+1 y=11 So this works out nice, your solutions for the set of equations is (5,11) since they are the same points

OpenStudy (kenkaneki12280000):

so its 5 and 11 right

OpenStudy (jameshorton):

yea

OpenStudy (kenkaneki12280000):

k ill tell if they are right

OpenStudy (kenkaneki12280000):

no i got a 40%

OpenStudy (jameshorton):

lol

OpenStudy (kenkaneki12280000):

we got 3 wrong

OpenStudy (jameshorton):

you in k12 or sum

OpenStudy (kenkaneki12280000):

yea